OpenStudy (anonymous):

a military tent of height 10m is in the form of a right circular cylinder with diameter of base 24m and height 5.00m surmounted by a right circular cone of same base radius. find the cost of canvas used in making the tent at rupess 42 per sq.m. (use piee = 22/7).

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Total height of the cone = 10 m Height of the cylindrical part = 5 m therefore height of conical part = 10m- 5m = 5m base radius of the cylindrical as well as conical part are same Hence r = 24/2= 12m for slant height of conical part, we have \[l^2= r^2+h^2 =12^2+5^2 = 144+25= 169\] \[\rightarrow l^2= 169 \rightarrow l = \sqrt {169} = 13 m\] Now, in order to find the canvas used in making the tent, we need to calculate Canvas used in making the tent= CSA of cylindrical part + CSA of conical part \[2 \pi r h + \pi r l = \pi r (2h+l) = \frac {22}{7} \times 12 ( 2 \times 5 + 13)\] \[= \frac {22}{7} \times 12 \times 23 = \frac {6072}{7}m^2\] Now, cost of canvas used in making the tent = \[\frac {6072}{7} m^2 \times 42 Rs/ m^2 =Rs. 36432 \]

OpenStudy (anonymous):

@meghanamaggi

OpenStudy (anonymous):

omg!!! thnku a lot, ...